# BoostStep: Boosting mathematical capability of Large Language Models via improved single-step reasoning

### v1

@@ Abstract
Cutting-edge large language models (LLMs) demonstrate promising performance in solving complex math problems with a divide-and-conquer pipeline and the assistance of in-context learning (ICL) examples. However, their potential for improvement is limited by two critical problems within their ICL examples: granularity-mismatch and the ensuing negative-effect noise problem. Specifically, the LLMs are capable of the dividing process yet mostly failed by inaccurate reasoning within a few conquer steps, while the ICL examples retrieved in question-grained sometimes lack relevant steps for a specific challenging reasoning step. Further, this disconnect may hinder the correct reasoning due to its irrelevance. To this end, we focus on improving the reasoning quality within each step and present BoostStep. BoostStep aligns the granularity between the retrieving and reasoning on step grained, and provides highly related ICL examples for each reasoning step with a novel ‘first-try’ strategy. BoostStep provides more relevant examples than the coarse question-grained strategy, enhancing the model reasoning quality within each step steadily. BoostStep is a general and robust reasoning-enhancing method that not only improves standalone reasoning performance but also integrates seamlessly with Monte Carlo Tree Search methods (MCTS) to refine both candidate generation and decision-making. Quantitatively, it improves GPT-4o and Qwen2.5-Math-72B by 3.6% and 2.0% respectively on various mathematical benchmarks, and 7.5% gain combined with MCTS.
@@ 1 Introduction
Figure 1: Our step-level in-context learning outperforms traditional problem-level few-shot learning for about 4% across in-domain, out-domain and cross-modality mathematical benchmark. Moreover, on benchmarks with lower similarity with example problem set (i.e. OlympiadBench and multi-modal benchmarks), where problem-level few-shot learning may have negative impact, while our strategy still provides valuable guidance.
Mathematical reasoning is a crucial and challenging task in the development of artificial intelligence. It serves as an indicator of a model’s ability to perform complex reasoning and has a wide range of applications, such as problem solving, theorem proving and scientific discovery.
@@
Through the analysis of error cases, we found that the current SOTA models have been relatively clear in the divide part, that is, the model can basically know exactly what tasks should be completed in each step. However, there will still be a lot of mistakes in the conquer section, such as wrong formula use, wrong calculation, insufficient enumeration, etc. To quantitatively substantiate this conclusion, we provide GPT-4o-mini with groundtruth reasoning process to determine whether the error in another response were due to an overarching flawed reasoning approach or a deviation within a particular step. In less advanced models like LLaMA-3.1-8B-Instruct Dubey et al. (2024), 8.7% of errors originates from an overall flawed reasoning approach. While in more advanced models like GPT-4o, only 0.8% of errors are ascribable to the overall approach, with the majority resulting from inaccuracies at individual steps of the reasoning process. Therefore, it can be concluded that the correctness of one-step reasoning is the bottleneck of reasoning capability.
Various approaches have been employed to improve reasoning correctness, such as producing chains of thought through prompt engineering Kojima et al. (2022); Wei et al. (2022), fine-tuning with mathematical data Shao et al. (2024); Yang et al. (2024); Ying et al. (2024), or generating multiple candidate reasoning paths using Monte Carlo Tree Search (MCTS) Zhang et al. (2024b, a).
Among those techniques, in-context learning is a particularly important one, which offers similar examples to provide detailed guidance. However, the examples retrieved by traditional problem-level in-context learning are listed before the reasoning process ever begins, thereby lacking fine-grained guidance during the reasoning process. Moreover, since the example problem can’t be exactly identical to the new one, the irrelevant steps in those examples may even become a distraction from the current reasoning, thus even negatively affecting the conquer capability for some specific steps.
To this end, we refine in-context learning from problem-level to step-level granularity to offer similar example steps during an on-going reasoning process for fine-grained step-level guidance, as well as ensuring that the introduced example is still highly relevant at the step level to avoid distractions.
@@
Building on the step-level granularity within the example problem bank, we propose an approach that incorporates in-context learning through a "first-try" format during an on-going reasoning process. Specifically, for a given problem to be solved, we break down the solving process into step-by-step reasoning paths. During the reasoning of a single step, we first allow the model to attempt a "first try" to comprehend what the model currently needs to reason about. Based on this initial attempt, we search the question bank to find similar steps that can guide the model to accurately output the current step. This first-try strategy helps ensure a higher similarity between the retrieved examples and the current step. Therefore, the distraction from irrelevant steps can be avoided and and the guidance effect is improved.
Compared with traditional in-context learning in problem-level granularity, our method is capable of providing examples during the reasoning process directly based on the steps to be solved, thereby offering more relevant guidance. Our approach demonstrates significant improvements over traditional few-shot learning across various open-source benchmarks, with an average increase of 4%.
@@
Our method significantly enhances single-step reasoning accuracy through step-level in-context learning, making it highly compatible with various current reasoning strategies that employ Monte Carlo Tree Search (MCTS). Typically, an MCTS method requires a reasoning model to generate multiple step-wise candidate reasoning paths and a critic model to evaluate the correctness of these candidates. Our approach can be integrated into both aspects in a plug-and-play manner. Specifically, when generating new candidate reasoning nodes, our method can introduce similar examples in the aforementioned ’first-try’ manner to improve the accuracy of candidates. Additionally, it can aid the critic model by incorporating similar correct reasoning steps into the evaluation of candidate reasoning processes. Experimental results indicate that both applications contribute positively to the model’s reasoning capabilities during MCTS process. Detailed illustration will be shown in Sec. 3.4
@@ 2 Related Works
Mathematical Reasoning. Mathematical reasoning has long been a highly challenging task in the field of artificial intelligence. In the early days of artificial intelligence, constrained by a lack of general capabilities, early methods Feigenbaum et al. (1963); Fletcher (1985) primarily attempted to perform simple mathematical reasoning through rule-based methods. With the advent of large language models with enhanced reasoning capabilities, contemporary approaches typically focus on enhancing performance during both the training and inference phases. The first category improves mathematical capability by fine-tuning with more high-quality mathematical data Shao et al. (2024); Yang et al. (2024); Lewkowycz et al. (2022); Yue et al. (2023); Xu et al. (2024). This strategy can fundamentally improve the base model’s mathematical capabilities. However, it demands substantial high-quality mathematical data and computational resources. Consequently, more efforts have been put in exploring various techniques during inference to enhance mathematical reasoning performance. Some work Wei et al. (2022); Kojima et al. (2022) involves prompt engineering to enable models to generate comprehensive chains of thought. Other studies Madaan et al. (2024); Gou et al. (2023); Ke et al. (2024) involve using self-refinement techniques to revise the initial reasoning outputs.
Stepwise Mathematical Reasoning. Recently, to further enhance mathematical reasoning capabilities, many studies have shifted the granularity of mathematical reasoning from the problem level to the step level. This approach involves addressing each next step individually, completing small segments of reasoning within the overall task. These works often employ strategies such as Tree of Thoughts (ToT) Yao et al. (2024) or Monte Carlo Tree Search Zhang et al. (2024b, a); Chen et al. (2024); Feng et al. (2023); Zhu et al. (2022), extending multiple steps to optimize step answers and ultimately obtain the optimal solution. Additionally, Process Supervision Models (PRMs) Lightman et al. (2023); Luo et al. (2024) are frequently used to verify the correctness of new candidate nodes in real-time and prune reasoning paths, thereby improving the accuracy of the final answer. This more detailed auxiliary strategy demonstrates greater potential.
@@ 3 Step-Level In-Context Learning
Figure 2: Our strategy refines in-context learning from problem-level granularity (fig.a) to step-level granularity(fig.b) to provide more real-time fine-grained guidance. Moreover, our strategy can guide the reasoning and verifying process in Monte Carlo Tree Search (MCTS) strategies by introducing examples.
@@ 3.1 Revisiting In-Context Learning from Conditional Probability
In-context learning provides the model with conditional probabilities similar to the groundtruth answer for imitation without changing the probability model $P_{model}$. Specifically, an example problem $q^{\prime}$ and its corresponding correct solution $r^{\prime}$ is provided and it can be posited that the conditional probability $P(r^{\prime}\mid q^{\prime})$ is similar to the probability of the grountruth answer of the target problem $P(r_{gt}\mid q)$. Consequently, the model will imitate this similar example and $r^{\prime}_{predict}=\arg\max\limits_{r}P_{model}(r\mid q,q^{\prime},r^{\prime})$ will be closer to $r_{gt}$ comparing to $r_{predict}$.
@@
To this end, we propose step-level in-context learning and first-try strategy to provide detailed and relevant example step when in step-level reasoning. Specifically, when generating new steps $s_{i+1}$ based on previous reasoning steps $s_{i},s_{i-1},\ldots,s_{1}$ and question $q$, we first utilize a first-try strategy to obtain an approximate estimate of $s^{first}_{i+1}$. Then, we use this $s^{first}_{i+1}$ to retrieve a similar step $s^{\prime}_{n+1}$ along with the corresponding $q^{\prime},s^{\prime}_{1},s^{\prime}_{2},\ldots,s^{\prime}_{n}$. Since these two steps are similar, a very reasonable assumption is that $P(s^{\prime}_{n+1}\mid q^{\prime},s^{\prime}_{1},\ldots,s^{\prime}_{n})$ closely approximates $P(s_{{gt}_{i+1}}\mid q,s_{1},\ldots,s_{i})$. Therefore, the generated step $s_{i+1}=\arg\max\limits_{s}P_{model}(s\mid q,s_{1},\ldots,s_{i},s^{\prime}_{n+1},s^{\prime}_{n},\ldots,s^{\prime}_{1},q^{\prime})$ will be more closed to $s_{{gt}_{i+1}}$ comparing to $s^{0-shot}_{i+1}$. Details about our step-level in-context learning and first-try strategy will be explained in Sec. 3.3
@@ 3.2 Step-Level Example Problem Bank
Figure 3: An example of different problems may contain similar steps. Problem-level in-context learning will ignore this example due to low problem similarity, while our step-level in-context learning can introduce the core skills by step-level retrieval and guidance.
Due to the need for further improvement in mathematical capabilities, currently open-source mathematical data no longer consist solely of problems and their final answers to determine whether the final answer obtained is correct or not. Instead, they also provide detailed solution processes to provide more fine-grained measurement. However, most current open-source mathematical data still do not break down the solution processes to the step level.
Some approaches Lightman et al. (2023) proposed using the period ’.’ as a clear semantic delimiter to segment steps for training PRM. This strategy allows for the quick decomposition of each step from a complete process without any additional assistance. However, this simple decomposition mode is obviously unreliable. Essentially, a single reasoning step should have a consistent target and encompass a complete thought process, making it the atomic granularity of reasoning. Using a period ’.’ as a delimiter may disrupt this atomicity. For example, it may split a complete enumerations for the same objective into multiple steps.
@@
The proposed step-level example problem bank is available at https://github.com/beichenzbc/BoostStep
@@ 3.3 Step-Level ICL with First-try Strategy
To this end, we propose a straightforward and effective "first-try" strategy to enhance the similarity of search steps. Our premise is that the most accurate way to estimate the next step is to actually allow the model to attempt the reasoning for the next step. Specifically, given a problem $q$ and all preceding reasoning steps $s_{i-1},s_{i-2},\ldots,s_{1}$, we first instruct the model to attempt continuing the reasoning process to arrive at a tentative step $s_{i}^{try}$ without the aid of any examples. Subsequently, we use $s_{i}^{try}$ to retrieve similar steps $s^{\prime}_{j}$ along with their corresponding problem $q^{\prime}$ and preceding steps $s^{\prime}_{1},\ldots,s^{\prime}_{j-1}$ from a step-level database. Finally, we feed the retrieved similar steps back to the model, enabling it to deduce the final step $s_{i}$. Besides, we add a widely-accepted strategy reference rejection. Specifically, If the similarity of the retrieved most similar example remains below a certain threshold, we consider that there are no sufficiently similar examples available for reference. Consequently, we do not provide any examples in order to avoid the negative effects associated with incoherent in-context learning. This "try-retrieve-reason" strategy significantly enhances retrieval relevance, thereby improving reasoning effectiveness. Experiments in Sec. 4.4 compare our method with several other retrieval strategies, demonstrating the superiority of our approach.
3.4 Step-Level Example Guidance in MCTS
The core aspect of our method is that our step-level in-context learning can significantly enhance the model’s single-step reasoning capability, which makes it easily integrable into common step-reasoning strategies, such as Monte Carlo Tree Search (MCTS).
Generally, MCTS methods necessitate two key components: a reasoning model that generates step-level reasoning and a Process-Supervision Reward Model (PRM) that continuously evaluates the current reasoning step in real time. Our method is beneficial for both of these components. It enhances the step-level reasoning performed by the reasoning model and improves the effectiveness of the PRM in evaluating current reasoning steps.
For the reasoning model, MCTS inherently requires step-by-step reasoning expansion. When expanding at node $s_{i}$, we simply apply the previously mentioned strategy: the model performs $n$ first tries to obtain $n$ example steps. For each example, the model completes the reasoning to generate $n$ child nodes $s^{1}_{i+1},s^{2}_{i+1},\ldots,s^{n}_{i+1}$. Our strategy improves the accuracy of individual child nodes $s^{j}_{i+1}$, thereby enhancing the overall effectiveness of the MCTS strategy.

### v2

@@ Abstract
Cutting-edge large language models (LLMs) demonstrate promising performance in solving complex math problems with a divide-and-conquer pipeline and the assistance of in-context learning (ICL) examples. However, their potential for improvement is limited by two critical problems within their ICL examples: granularity-mismatch and the ensuing negative-effect noise problem. Specifically, the LLMs are capable of the dividing process yet mostly fail by inaccurate reasoning within a few conquer steps, while the ICL examples retrieved in question-grained sometimes lack relevant steps for a specific challenging reasoning step. Further, this disconnect may hinder the correct reasoning due to its irrelevance. To this end, we focus on improving the reasoning quality within each step and present BoostStep. BoostStep aligns the granularity between the retrieving and reasoning on step-grained and provides highly related ICL examples for each reasoning step with a novel ‘first-try’ strategy. BoostStep provides more relevant examples than the coarse question-grained strategy, enhancing the model reasoning quality within each step steadily. BoostStep is a general and robust reasoning-enhancing method that not only improves standalone reasoning performance but also integrates seamlessly with Step-level tree search methods to refine both candidate generation and decision-making. Quantitatively, it improves GPT-4o and Qwen2.5-Math-72B by 3.6% and 2.0% respectively on various mathematical benchmarks, and 7.5% gain combined with tree search.
@@ 1 Introduction
Figure 1: Our step-level in-context learning outperforms traditional problem-level few-shot learning for about 4% across in-domain, out-domain and cross-modality mathematical benchmark. Moreover, on benchmarks with lower similarity with the example problem set (i.e. OlympiadBench and multi-modal benchmarks), where problem-level few-shot learning may have a negative impact, our strategy still provides valuable guidance.
Mathematical reasoning is a crucial and challenging task in the development of artificial intelligence. It serves as an indicator of a model’s ability to perform complex reasoning and has a wide range of applications, such as problem-solving, theorem proving, and scientific discovery.
@@
Through the analysis of error cases, we found that the current SOTA models have been relatively clear in the divide part, that is, the model can know exactly what tasks should be completed in each step. However, there will still be a lot of mistakes in the conquer section, such as wrong formula use, wrong calculation, insufficient enumeration, etc. To quantitatively substantiate this conclusion, we provide GPT-4o-mini with a ground truth reasoning process to determine whether the error in another response was due to an overarching flawed reasoning approach or a deviation within a particular step. In less advanced models like LLaMA-3.1-8B-Instruct Dubey et al. (2024), 8.7% of errors originate from an overall flawed reasoning approach. In more advanced models like GPT-4o, only 0.8% of errors are ascribable to the overall approach, with the majority resulting from inaccuracies at individual steps of the reasoning process. Therefore, the correctness of one-step reasoning is the bottleneck of reasoning capability.
Various approaches have been employed to improve reasoning correctness, such as producing chains of thought through prompt engineering Kojima et al. (2022); Wei et al. (2022), fine-tuning with mathematical data Shao et al. (2024); Yang et al. (2024); Ying et al. (2024), or generating multiple candidate reasoning paths using Tree Search Methods Zhang et al. (2024b, a).
Among those techniques, in-context learning is a particularly important one, which offers similar examples to provide detailed guidance. However, the examples retrieved by traditional problem-level in-context learning are listed before the reasoning process ever begins, thereby lacking fine-grained guidance during the reasoning process. Moreover, since the example problem can’t be identical to the new one, the irrelevant steps in those examples may even become a distraction from the current reasoning, thus even negatively affecting the conquer capability for some specific steps.
To this end, we refine in-context learning from problem-level to step-level granularity to offer similar example steps during an ongoing reasoning process for fine-grained step-level guidance, as well as ensuring that the introduced example is still highly relevant at the step level to avoid distractions.
@@
Building on the step-level granularity within the example problem bank, we propose an approach that incorporates in-context learning through a "first-try" format during an ongoing reasoning process. Specifically, for a given problem to be solved, we break down the solving process into step-by-step reasoning paths. During the reasoning of a single step, we first allow the model to attempt a ‘first try’ to comprehend what the model currently needs to reason about. Based on this initial attempt, we searched the question bank to find similar steps that can guide the model to accurately output the current step. This first-try strategy helps ensure a higher similarity between the retrieved examples and the current step. Therefore, the distraction from irrelevant steps can be avoided and the guidance effect is improved.
Compared with traditional in-context learning in problem-level granularity, our method is capable of providing examples during the reasoning process directly based on the steps to be solved, thereby offering more relevant guidance. Our approach demonstrates significant improvements over traditional few-shot learning across various open-source benchmarks, with an average increase of 3.6%.
@@
Our method significantly enhances single-step reasoning accuracy through step-level in-context learning, making it highly compatible with various current reasoning strategies that employ step-level tree search. Typically, a tree-search method requires a reasoning model to generate multiple step-wise candidate reasoning paths and a critic model to evaluate the correctness of these candidates. Our approach can be integrated into both aspects in a plug-and-play manner. Specifically, when the reasoning model generates new candidate reasoning nodes, our method can introduce similar examples in the aforementioned ‘first-try’ manner to improve the accuracy of candidates. Additionally, it can aid the critic model by incorporating similar correct reasoning steps into the evaluation of candidate reasoning processes to provide similar guidance. Experimental results indicate that both applications contribute positively to the model’s reasoning capabilities. Detailed illustration will be shown in Sec. 3.4
@@ 2 Related Works
Mathematical Reasoning. Mathematical reasoning has long been a highly challenging task in the field of artificial intelligence. In the early days of artificial intelligence, constrained by a lack of general capabilities, early methods Feigenbaum et al. (1963); Fletcher (1985) primarily attempted to perform simple mathematical reasoning through rule-based methods. With the advent of large language models with enhanced reasoning capabilities, contemporary approaches typically focus on enhancing performance during both the training and inference phases. The first category improves mathematical capability by fine-tuning with more high-quality mathematical data Shao et al. (2024); Yang et al. (2024); Lewkowycz et al. (2022); Yue et al. (2023); Xu et al. (2024). This strategy can fundamentally improve the base model’s mathematical capabilities. However, it demands substantial high-quality mathematical data and computational resources. Consequently, more efforts have been put into exploring various techniques during inference to enhance mathematical reasoning performance. Some work Wei et al. (2022); Kojima et al. (2022) involves prompt engineering to enable models to generate comprehensive chains of thought. Other studies Madaan et al. (2024); Gou et al. (2023); Ke et al. (2024) use self-refinement techniques to revise the initial reasoning outputs.
Stepwise Mathematical Reasoning. Recently, to further enhance mathematical reasoning capabilities, many studies have shifted the granularity of mathematical reasoning from the problem level to the step level. This approach involves addressing each next step individually and completing small segments of reasoning within the overall task. These works often employ tree searching strategies like Tree of Thoughts (ToT) Yao et al. (2024) or Monte Carlo Tree Search Zhang et al. (2024b, a); Chen et al. (2024); Feng et al. (2023); Zhu et al. (2022), extending multiple steps to optimize step answers and ultimately obtain the optimal solution. Additionally, Process Supervision Models (PRMs) Lightman et al. (2023); Luo et al. (2024) are frequently used to verify the correctness of new candidate nodes in real-time and prune reasoning paths, thereby improving the accuracy of the final answer. This more detailed auxiliary strategy demonstrates greater potential.
@@ 3 Step-Level In-Context Learning
Figure 2: Our strategy refines in-context learning from problem-level granularity (fig.a) to step-level granularity(fig.b) to provide more real-time fine-grained guidance. Moreover, our strategy can guide the reasoning and verifying process in tree-searching strategies by introducing examples.
@@ 3.1 Revisiting In-Context Learning from Conditional Probability
In-context learning provides the model with conditional probabilities similar to the ground truth answer for imitation without changing the probability model $P_{model}$. Specifically, an example problem $q^{\prime}$ and its corresponding correct solution $r^{\prime}$ is provided and it can be posited that the conditional probability $P(r^{\prime}\mid q^{\prime})$ is similar to the probability of the ground truth answer of the target problem $P(r_{gt}\mid q)$. Consequently, the model will imitate this similar example and $r^{\prime}_{predict}=\arg\max\limits_{r}P_{model}(r\mid q,q^{\prime},r^{\prime})$ will be closer to $r_{gt}$ comparing to $r_{predict}$.
@@
To this end, we propose step-level in-context learning and a first-try strategy to provide detailed and relevant example steps when in step-level reasoning. Specifically, when generating new steps $s_{i+1}$ based on previous reasoning steps $s_{i},s_{i-1},\ldots,s_{1}$ and question $q$, we first utilize a first-try strategy to obtain an approximate estimate of $s^{first}_{i+1}$. Then, we use this $s^{first}_{i+1}$ to retrieve a similar step $s^{\prime}_{n+1}$ along with the corresponding $q^{\prime},s^{\prime}_{1},s^{\prime}_{2},\ldots,s^{\prime}_{n}$. Since these two steps are similar, a very reasonable assumption is that $P(s^{\prime}_{n+1}\mid q^{\prime},s^{\prime}_{1},\ldots,s^{\prime}_{n})$ closely approximates $P(s_{{gt}_{i+1}}\mid q,s_{1},\ldots,s_{i})$. Therefore, the generated step $s_{i+1}=\arg\max\limits_{s}P_{model}(s\mid q,s_{1},\ldots,s_{i},q^{\prime},s^{\prime}_{1},\ldots,s^{\prime}_{n},,s^{\prime}_{n+1},)$ will be more closed to $s_{{gt}_{i+1}}$ comparing to $s^{0-shot}_{i+1}$. Details about our step-level in-context learning and first-try strategy will be explained in Sec. 3.3
@@ 3.2 Step-Level Example Problem Bank
Figure 3: Different problems may contain similar steps. Problem-level in-context learning will ignore this example due to low problem similarity. In contrast, our step-level in-context learning strategy can introduce the core skills by step-level retrieval and guidance.
Due to the need for further improvement in mathematical capabilities, current open-source mathematical data no longer consist solely of problems and their final answers to determine whether the final answer obtained is correct or not. Instead, they also provide detailed solution processes to provide more fine-grained measurements. However, most current open-source mathematical data still do not break down the solution processes to the step level.
Some approaches Lightman et al. (2023) proposed using the period ’.’ as a clear semantic delimiter to segment steps for training PRM. This strategy allows for the quick decomposition of each step from a complete process without any additional assistance. However, this simple decomposition mode is obviously unreliable. Essentially, a single reasoning step should have a consistent target and encompass a complete thought process, making it the atomic granularity of reasoning. Using a period ’.’ as a delimiter may disrupt this atomicity. For example, it may split a complete enumeration for the same objective into multiple steps.
@@
* The proposed step-level example problem bank is available at https://github.com/beichenzbc/BoostStep
@@ 3.3 Step-Level ICL with First-try Strategy
To this end, we propose a straightforward and effective "first-try" strategy to enhance the similarity of search steps. Our premise is that the most accurate way to estimate the next step is to actually allow the model to attempt the reasoning for the next step. Specifically, given a problem $q$ and all preceding reasoning steps $s_{i-1},s_{i-2},\ldots,s_{1}$, we first instruct the model to attempt continuing the reasoning process to arrive at a tentative step $s_{i}^{try}$ without the aid of any examples. Subsequently, we use $s_{i}^{try}$ to retrieve similar steps $s^{\prime}_{j}$ along with their corresponding problem $q^{\prime}$ and preceding steps $s^{\prime}_{1},\ldots,s^{\prime}_{j-1}$ from a step-level database. Finally, we feed the retrieved similar steps back to the model, enabling it to deduce the final step $s_{i}$. Besides, we add a widely accepted strategy reference rejection. Specifically, if the similarity of the retrieved most similar example remains below a certain threshold, we consider that there are no sufficiently similar examples available for reference. Consequently, we do not provide any examples to avoid the negative effects associated with incoherent in-context learning. This "try-retrieve-reason" strategy significantly enhances retrieval relevance, thereby improving reasoning effectiveness. Experiments in Sec. 4.4 compare our method with several other retrieval strategies, demonstrating the superiority of our approach.
3.4 Step-Level Guidance in Tree Search
Our step-level in-context learning can significantly enhance the model’s single-step reasoning capability, which makes it easily integrated into common step-level tree-search strategies.
Generally, tree search methods necessitate two key components: a reasoning model that generates step-level reasoning and a Process-Supervision Reward Model (PRM) that continuously evaluates the current reasoning step in real time. Our method is beneficial for both of these components. It enhances the step-level reasoning performed by the reasoning model and improves the effectiveness of the PRM in evaluating current reasoning steps.
For the reasoning model, tree search methods inherently require step-by-step reasoning expansion. When expanding at node $s_{i}$, we simply apply the previously mentioned strategy: the model performs $n$ first tries to obtain $n$ example steps. For each example, the model then completes the reasoning to generate $n$ child nodes $s^{1}_{i+1},s^{2}_{i+1},\ldots,s^{n}_{i+1}$ with the help of these examples. Our strategy improves the accuracy of individual child nodes $s^{j}_{i+1}$, thereby enhancing the overall reasoning quality.